Student Expenditures The average expenditure per student (based on average daily attendance) for a certain school year was $10,337 with a population standard deviation of $1560. A survey for the next school year of 150 randomly selected students resulted in a sample mean of $10,798. Do these results indicate that the average expenditure has changed? Choose your own level of significance.


1
Expert's answer
2022-01-04T11:28:43-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=10337"

"H_1:\mu\not=10337"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{10798−10337}{1560/\sqrt{150}}"

"\approx3.6193"

Since it is observed that "|z| = 3.6193 >1.96= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z>3.6193)\approx 0.000295," and since "p= 0.000295<0.05=\alpha," 0.05 it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "10337," at the "\alpha = 0.05" significance level.

Therefore, there is enough evidence to claim that the average expenditure has changed, at the "\alpha = 0.05" significance level.